40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 27 16:28:57 2022, Anchal, Summary, BHD, BHD Upgrade - Retreiving arm cavity alignment 
    Reply  Thu Apr 28 09:12:18 2022, Anchal, Update, BHD, Restoring arm alignment 
       Reply  Thu Apr 28 18:07:04 2022, Anchal, Update, BHD, Restoring arm alignment at EOD 
Message ID: 16816     Entry time: Thu Apr 28 09:12:18 2022     In reply to: 16815     Reply to this: 16819
Author: Anchal 
Type: Update 
Category: BHD 
Subject: Restoring arm alignment 

This would be a daily first task in the morning. We'll need to check the status of arm alignment and optimize it back to maximum every morning for the rest of the day's work.

Today, when I came, on openin gthe PSL shutter, IMC was aligned good, both arms were flashing but YARM maximum transmission count was around 0.7 (as opposed to 1 from yesterday) and XARM maximum transmission count was 0.5 (as opposed to 1 from yesterday). I did not change the input alignment to the interferometer. I only used ITMY-ETMY to regain flashing count of 1 on YARM and used BS and tehn ITMX-ETMX to regain flashing count of 0.9 to 1 in XARM.

Even thought the oplevs were centered yesterday, I found the oplev had drifted from the center and the optimal position also is different for all ooptics except EMTY and BS. It is worth nothign that in optimal position both PIT and YAW of ITMY and ITMX are off by 70-90 uradians and ETMX Pit oplev is off by 55 uradians.

 

ELOG V3.1.3-